Without using software, please complete the following question by hand:

(3) Sketch the solid over which the iterated integral LL*65(71.4,2) de dr dy is evaluated.
Answer #1

Y+3 x+4 let IF f(xiy, 2) dzdxdy 4 x+2 2 = X+2 to 2= x +4 X- y to x = y + 3 ý=1 y = 1 Z 10 8 6 2 -6 -2 oplo 2 2 -2 y -4 -6 -8